# Late Chunking: Contextual Chunk Embeddings Using Long-Context Embedding Models

### v2

@@ 1 Introduction
Neural information retrieval (IR) relies on text embedding models (Reimers & Gurevych, 2019) that are primarily based on the transformer architecture (Devlin et al., 2019) and have been pre-trained using very large text corpora. These models capture important elements of texts’ semantics in the form of dense vectors whose spatial relations - particularly cosine distance - are good proxies for text similarity and relevancy. For many neural IR use cases like the well-known RAG (Retrieval Augmented Generation) approach (Lewis et al., 2020), applications require splitting documents into limited-size text chunks, and storing them and their vector embeddings in a database. At run-time, neural IR techniques are used to retrieve chunks of text relevant to a user’s requests, which are, in the case of RAG, presented to an LLM as a basis for synthesizing a response.
@@
To overcome this limitation, we introduce a novel technique called late chunking. This method leverages the long text embedding capabilities of recently published open source models (Günther et al., 2023; Nussbaum et al., 2024) to, first, encode all tokens of an entire document with their full in-document context into a sequence of token embeddings, and then break this sequence up into chunks, which receive embeddings via mean pooling of their token embeddings. This way, chunk embeddings include relevant semantic information derived from their place in the whole text.
@@ 2 Related Work
To prevent the problem of missing context information various approaches have been proposed that augment the text of the chunks. For instance, practitioners divide text into overlapping chunks (Safjan, 2023), meaning that the end of one chunk shares some tokens with the beginning of the next chunk. During the development of this paper, a blog post (Anthropic, 2024) introduced an alternative approach for producing contextualized chunk embeddings using an additional large language model (LLM). The LLM receives as input the whole document and the target chunk to produce text for augmenting the chunk text with relevant context information before passing it to the embedding model. This is however computationally more expensive, as LLMs are typically much larger than embedding models or even require paid access to LLM APIs.
@@ 3.2 Training Method
The fine-tuning procedure itself follows the pair training stage described in Günther et al. (2023), where the model is trained on text pairs using the InfoNCE (van den Oord et al., 2018) loss which is defined on a batch of pairs and the cosine similarity function :
@@ 4.1 Evaluation on Retrieval Tasks
We present the evaluation results in Table 2. When comparing the results for the different chunking methods, we observe that replacing naive methods with their late chunking counterparts almost always yields better performance. Averaging results across three models and four datasets, we find a 3.63% relative improvement (1.9% absolute) from naive chunking with sentence boundaries to late chunking using sentence boundaries, a 3.46% improvement (1.8% absolute) from naive chunking to late chunking using fixed-size boundaries, and a 2.70% improvement (1.5% absolute) from naive chunking to late chunking when using semantic sentence boundaries. These findings demonstrate that the late chunking technique effectively and consistently enhances overall performance.
@@ 4.4 Evaluation of Training Method
For the span-based training method, we prepare two datasets into the format described in Section 3.2 and make these publicly available on HuggingFace 444https://huggingface.co/datasets/jinaai/fever-span-annotated; https://huggingface.co/datasets/jinaai/triviaqa-span-annotated . These two datasets are FEVER (Thorne et al., 2018) and TriviaQA (Joshi et al., 2017), which are well-suited for this experiment as they contain annotations of where the relevant text can be found in the documents respectively. In the FEVER dataset, these spans take the shape of sentence number annotations, while for TriviaQA the annotations are usually a name, place, or date in the form of a short phrase. For FEVER, we only include pairs where the document provides supporting evidence for the claim. When multiple spans are annotated in these datasets, we select only the span, which occurs earliest in the document.

### v3

@@ 1 Introduction
Neural information retrieval (IR) relies on text embedding models (Reimers & Gurevych, 2019) that are primarily based on the transformer architecture (Devlin et al., 2019) and have been pre-trained using very large text corpora. These models capture important elements of texts’ semantics in the form of dense vectors whose spatial relations - particularly cosine distance - are good proxies for text similarity and relevancy. For many neural IR use cases like the well-known RAG (Retrieval Augmented Generation) approach (Lewis et al., 2020), applications require splitting documents into limited-size text chunks, and storing them and their vector embeddings in a database. At run-time, neural IR techniques are used to retrieve chunks of text relevant to a user’s requests, which are, in the case of RAG, presented to an LLM as a basis for synthesizing a response. Furthermore, many other applications require processing small text segments (Salton et al., 1993; Demszky et al., 2020), and therefore rely on chunking, e.g., to quickly navigate a user to the relevant passage in a document Callan (1994).
Moreover, while long context embedding models can improve retrieval performance on long texts (Günther et al., 2023), they still perform better on short texts (Zhou et al., 2024). As a result, chunking generally improves retrieval, even with models that support long contexts. (See Appendix A.1.)
@@
To overcome this limitation, we introduce a novel technique called late chunking. This method leverages the long text embedding capabilities of recently published models (Günther et al., 2023; Nussbaum et al., 2024) to, first, encode all tokens of an entire document with their full in-document context into a sequence of token embeddings, and then break this sequence up into chunks, which receive embeddings via mean pooling of their token embeddings. This way, chunk embeddings include relevant semantic information derived from their place in the whole text.
@@ 2 Related Work
To prevent the problem of missing context information various approaches have been proposed that augment the text of the chunks. For instance, practitioners divide text into overlapping chunks (Safjan, 2023), meaning that the end of one chunk shares some tokens with the beginning of the next chunk. During the development of this paper, a blog post (Anthropic, 2024) introduced an alternative approach for producing contextualized chunk embeddings using an additional large language model (LLM). The LLM receives as input the whole document and the target chunk to produce text for augmenting the chunk text with relevant context information before passing it to the embedding model. This is however computationally more expensive, as LLMs are typically much larger than embedding models or even require paid access to LLM APIs. Similarly, Luo et al. (2024) extract propositions for each paragraph using an additional language model. However, each paragraph is processed independently, which might result in losing context across paragraphs. Moreover, this approach cannot be used with any technique for segmenting the text (e.g. fixed-size, sentence-based, semantic chunking, …) but is restricted to the texts produced by the language model.
@@
Furthermore, Chen et al. (2024) aims to produce contextualized embedding representations by training an embedding model specifically to produce contextualized embedding representations of sentences.
@@ 3.2 Training Method
The fine-tuning procedure itself follows the pair training stage described in Günther et al. (2023), where the model is trained on text pairs using the InfoNCE loss function (van den Oord et al., 2018) which is defined on a batch of pairs and the cosine similarity function :
@@ 4.1 Evaluation on Retrieval Tasks
We present the evaluation results in Table 2. When comparing the results for the different chunking methods, we observe that replacing naive methods with their late chunking counterparts almost always yields better performance. Averaging results across three models and four datasets, we find a 3.63% relative improvement (1.9% absolute) from naive chunking with sentence boundaries to late chunking using sentence boundaries, a 3.46% improvement (1.8% absolute) from naive chunking to late chunking using fixed-size boundaries, and a 2.70% improvement (1.5% absolute) from naive chunking to late chunking when using semantic sentence boundaries. In all experiments the chunks are non-overlapping, however, additional results demonstrated in appendix A.2 show that overlapping the chunks generally neither improves nor harms the retrieval performance. These findings demonstrate that the late chunking technique effectively and consistently enhances overall performance.
@@ 4.4 Evaluation of Training Method
For the span-based training method, we prepare two datasets into the format described in Section 3.2 and make these publicly available on HuggingFace 444FEVER: https://huggingface.co/datasets/jinaai/fever-span-annotated
Trivia-QA: https://huggingface.co/datasets/jinaai/triviaqa-span-annotated. These two datasets are FEVER (Thorne et al., 2018) and TriviaQA (Joshi et al., 2017), which are well-suited for this experiment as they contain annotations of where the relevant text can be found in the documents respectively. In the FEVER dataset, these spans take the shape of sentence number annotations, while for TriviaQA the annotations are usually a name, place, or date in the form of a short phrase. For FEVER, we only include pairs where the document provides supporting evidence for the claim. When multiple spans are annotated in these datasets, we select only the span, which occurs earliest in the document.
@@ 5 Conclusion
Appendix A Appendix
A.1 Limitations of Long-Context Embedding Models
Model: jina-embeddings-v2-small, naive (not late) chunking with fixed-size strategy.
| NarrativeQA Dataset | |||
| Max. Length (Trucation) | Chunk-Size | Chunking | nDCG@10 |
| 192 | 192 | 20.26 | |
| 8192 | 8192 | 32.73 | |
| 8192 | 128 | ✓ | 46.28 |
| 8192 | 512 | ✓ | 47.63 |
| 2WikiMultiHopQA Dataset | |||
| Max. Length (Trucation) | Chunk-Size | Chunking | nDCG@10 |
| 192 | 192 | 48.86 | |
| 8192 | 8192 | 70.32 | |
| 8192 | 128 | ✓ | 91.36 |
| 8192 | 512 | ✓ | 86.3 |
| SummScreenFD Dataset | |||
| Max. Length (Trucation) | Chunk-Size | Chunking | nDCG@10 |
| 192 | 192 | 52.89 | |
| 8192 | 8192 | 91.24 | |
| 8192 | 128 | ✓ | 88.21 |
| 8192 | 512 | ✓ | 89.71 |
| QMSum Dataset | |||
| Max. Length (Trucation) | Chunk-Size | Chunking | nDCG@10 |
| 192 | 192 | 14.45 | |
| 8192 | 8192 | 36.81 | |
| 8192 | 128 | ✓ | 47.99 |
| 8192 | 512 | ✓ | 48.34 |
To investigate whether chunking is beneficial for retrieval tasks when all texts are smaller than the maximum input token length of the model, we truncate texts to a fixed length and apply chunking afterwards. We then evaluate the retrieval performance with the same setup as in Section 4.1, using the jina-embeddings-v2-small model and applying all the non-synthetic retrieval tasks from the LongEmbed benchmark (Zhu et al., 2024). We use fixed-size chunking (see Section 4.1) with chunk sizes of 128 and 512 tokens. Table 5 demonstrates that retrieval with chunking significantly outperforms retrieval without chunking. The average relative improvement from chunking with 512 tokens is +24.47%. Only on the SummScreenFD task did retrieval without chunking perform slightly better. Furthermore, truncating at 8192 tokens generally performs better than truncating at 192 tokens, indicating that long-text embedding models still provide an advantage over embedding models with short context lengths.
A.2 Retrieval with Overlapping Chunks
Model: jina-embeddings-v2-small. the fixed-size strategy (256 tokens), optional overlap (16 tokens). Scores are in nDCG@10 [%]
| Dataset | Naive Chunking | Late Chunking | ||
| w/ Overlap | w/o Overlap | w/ Overlap | w/o Overlap | |
| SciFact | 64.2 | 61.7 | 66.1 | 65.9 |
| NFCorpus | 23.5 | 22.8 | 30.0 | 30.5 |
| FiQA | 33.3 | 32.8 | 33.8 | 34.0 |
| TRECCOVID | 63.4 | 64.5 | 64.7 | 64.9 |
In practice, engineers often construct overlapping chunks to prevent the loss of context at the chunk boundaries (Safjan, 2023). To analyze whether this improves retrieval performance with jina-embeddings-v2-small and the evaluation setup from Section 4.1, we ran the BeIR benchmark tasks using fixed-size chunking (256 tokens) both with and without an overlap of 16 tokens. The results in Table 6 do not show a clear advantage of using overlaps. The nDCG@10 scores are relatively similar regardless.